### v1

@@
Abstract
3D Gaussian Splatting (3DGS) has been widely used in 3D reconstruction and 3D generation. Training to get a 3DGS scene often takes a lot of time and resources and even valuable inspiration. The increasing amount of 3DGS digital asset have brought great challenges to the copyright protection. However, it still lacks profound exploration targeted at 3DGS. In this paper, we propose a new framework X-SG2S which can simultaneously watermark 1 to 3D messages while keeping the original 3DGS scene almost unchanged. Generally, we have a X-SG2S injector for adding multi-modal messages simultaneously and an extractor for extract them. Specifically, we first split the watermarks into message patches in a fixed manner and sort the 3DGS points. A self-adaption gate is used to pick out suitable location for watermarking. Then use a XD(multi-dimension)-injection heads to add multi-modal messages into sorted 3DGS points. A learnable gate can recognize the location with extra messages and XD-extraction heads can restore hidden messages from the location recommended by the learnable gate. Extensive experiments demonstrated that the proposed X-SG2S can effectively conceal multi modal messages without changing pretrained 3DGS pipeline or the original form of 3DGS parameters. Meanwhile, with simple and efficient model structure and high practicality, X-SG2S still shows good performance in hiding and extracting multi-modal inner structured or unstructured messages. X-SG2S is the first to unify 1 to 3D watermarking model for 3DGS and the first framework to add multi-modal watermarks simultaneous in one 3DGS which pave the wave for later researches. Our project is available at: https://github.com/ChengLiDuoJi/XSGS.
ftcldj@mail.scut.edu.cn, ming.li@u.nus.edu, hpzhuang@scut.edu.cn
1 Introduction
Figure 1: Application scenario of the proposed. Scenario 1: after training a 3DGS pipeline by a trainer, an X-SG2S injector should be combined for joint use. In this way, when other users employ the trainer’s model, they can obtain a 3DGS result with some watermarks. By providing the 3DGS file to X-SG2S’s extractor, the trainer can determine whether a user has generated the result by using the trainer’s model. Scenario 2: if a user has specific requirements to add extra information to the 3DGS result for a particular purpose, they can directly use X-SG2S to achieve this goal. For example, you can add a text with “ICML”, an image of the ICML logo and a small 3DGS object at the same time by using X-SG2S’s injector and extract them by using X-SG2S’s extractor.
@@ 1 Introduction
Currently, only a small portion of research has addressed these problems. A new study GS-Hider restore image or 3D scene into another 3D scene via design a coupled secured feature. Considering the security issues of generative 3D, GaussianStego explores how to embed image watermarks into diffusion models. However, these models all have their limitations. For example, GS-Hider modifies the parameters of GS, which may increase the size of the 3D model parameters and no longer make it universal. The approach of GaussianStego rooting dino into diffusion can not well demonstrate its effect in feed forward 3D generative models. At the same time, adding extra elements in original 3DGS pipelines for protecting copyright is costly and complicate in design or training. We can see that how to effectively add additional watermarks or information without changing the GS parameters or model structure and let the method be general will be a great challenge. Meanwhile, embedding multimodal watermarks into a 3DGS simultaneously is even more of an unexplored area.
To solve the above challenges, we propose an effective, simple and flexible watermarking framework X-SG2S. It aims to embed 3D objects or images or binary messages simultaneously into the original 3DGS scene, and accurately extract the hidden message via just an ultra simple way. Specifically, we use a determined sorting method to sort the GS point set, then use a self-adaption gate which can learn the adding position through the interaction between the additional information(watermarks) and the container(3DGS scenes). A XD-injection heads are used to add different modal messages respectively. Then a learnable gate is used to recognize the locations where we have added the messages. Finally, XD-extraction heads are used to restore the message added. The training for X-SG2S does not need to fine tune or redesign the pipeline. This property makes this model suitable not only for the feedforward model, but also for the model like diffusion which needs multiple iterations. And if you do not need to add additional information, you can use the original model in the fastest time without further fine-tuning. However, it is also possible to fine tune with the pipeline. In a nutshell, the contributions and advantages of our X-SG2S can be summarized as follows.
@@
2 Related Works
2.1 Generalizable GS
3D Gaussian Splatting (3DGS) Kerbl et al. (2023) has emerged as a powerful method for reconstructing and representing 3D scenes using millions of 3D Gaussians. There has been a significant amount of work that has achieved good results in the reconstruction of small objects Szymanowicz et al. (2024) Boss et al. (2024) and scenes Chen et al. (2025) Charatan et al. (2024). Additionally, some studies have utilized diffusion models to realize text-to-3D generation Li et al. (2024b), which also achieves promising outcomes.
2.2 3D Steganography
Steganography has been evolving over the decades. Some research has made 3D steganography achieve good results in explicit geometry like meshes and point clouds Ohbuchi et al. (2002) Zhu et al. (2024) Ferreira & Lima (2020) or implicit geometer like nerf Li et al. (2023)Luo et al. (2023). 3DGS is a new technology to represent explicit geometry. However, few works have been researched in steganography for 3DGS. GS-Hider Zhang et al. (2024)design a coupled secured feature attribute to replace the original 3DGS’s spherical harmonics coefficients and then use a scene decoder and a message decoder to disentangle the original RGB scene and the hidden message. GaussianStegoLi et al. (2024a) use dino to add image watermarks and use U-Net to extract them via multi rendered views. But there is not research about watermarking 1 to 3D messages to single 3DGS scene simultaneous while keeping GS parameters or model structure unchanged.
3 Safe Generalizable GS with X-dimensional watermarks
3.1 Preliminaries of Gaussian Splatting
@@ 3.2 Task Settings
3.3 Deep Thinking About Watermarking 3DGS
@@
  • If we want to know how to extract the information from the point set, we should firstly know which of them are added. This means you would better know the location of each point and the restoration order. In another words, you should make the point set an ordered set and a known order of watermark composition.
3.4 Overview
Figure 2: This figure shows the structure of X-SG2S. X-SG2S is made up of a watermark injector and a watermark extractor. The injector includes a self-adaption gate to propose the locations for adding message patches and three injection heads to add 1 to 3D message patches to GS points respectively. The extractor includes a learnable gate to recognize the locations added extra message and three extraction heads to extract 1 to 3D message patches respectively. Self-adaption gate is composed by an embedding layer, an interaction block and a top-k. Embedding layer is used to embed multi modal message patches to a general space and embed GS points to “gs-emb” space. Interaction block uses the embedded message to generate the final-score which will be a reference for adding locations. In the experiment part, we compare five different interaction blocks and choose the one shown in this figure. Via top-k, the recommended locations can be made as binary mask. Every time, only one modality’s adding locations are recommended. By multiple consecutive calculations, it will recommend the locations for all the modalities. The mask of one modality may be reused to generate masks for latter modalities, which prevents overlapping of adding locations.
@@ 3.4 Overview
3.5 Point Set Sorting
As can be seen above, before injecting information and extracting information, we should use the sorting method to sort the point set. If you want every time to get the same order for a point set, you should use a deterministic sorting method. There are many methods for sorting. Here we choose Hilbert filling curve to complete this task. This is because Hilbert filling curve is a kind of deterministic curve. Hilbert curves can map high-dimensional data to a one-dimensional space to get the order of high-dimensional data. Meanwhile, when the size of the data set increases, data access time and computation time are often affected by spatial locality. Hilbert filling curve can reduce unnecessary jumps in storage and retrieval and improve the processing efficiency of large-scale data sets through space filling. At the same time, the Hilbert filling curve can be accelerated by GPU, which can effectively reduce the operation time.
3.6 Self-Adaption Gate
@@
This component is to find out the points which are suitable for injecting the messages while ensuring that the original scenario is not disturbed so much. A simple idea is that let a gate to learn a score to measure the locations. The gate will give every point a comprehensive score. A higher score indicates that the location is more suitable for adding information.
@@
The comprehensive score is calculated from two scores:“self-score” and “cross-score”. “Self-score” is used to measure which points can significantly impact the 3DGS scenario and which are trivial. It is obvious that if we choose the trivial points to inject messages, we will not disturb the scenario too much. But only use “self-score” is not enough. We need to know where are more suitable for message to inject and extract. “Cross-score” is used to manage this problem.
@@
After we get two scores, we simply multiply two scores together to get the final score. By using top-k, we can set the recommended k positions to 1 and others to 0 to get the selection mask.
3.6.1 message embedding
@@
Different modalities belong to different vector spaces. If we want to calculate the interactions results uniformly and efficiently, we need to map them into a unified space first. Here we simply use three different linear layer to map 1,2,3D message to a general space. And for gs cloud, we also use a linear layer to map it to another high dimension space we call “gs-emb” space. The dimension of two space can be different.
3.6.2 induced set attention to produce self-score
@@
After we get embedded message, we can measure the interact scores. Attention mechanism is a very efficient method for interact different information. We use transformer-based model to calculate the “self-score”. But we find that if we simply use scale dot product attention, the memory will overflow because the number of GS points are large. However, to calculate “self-score” is merely want to measure the importance of individual information in the overall information instead of learning detailed information between individuals. So set attentionLee et al. (2019) is used to overcome the memory problem. At the same time, set transformer does not need for the data to be sequential. This is very suitable for point sets, because point sets are unordered.
@@
3.6.3 efficient attention to produce cross-score
Because “Cross-score” needs to interact the whole GS points with embedded message patches. However, if the amount of message patches is large, the attention may also face the problem of overflowing. Thus, we use efficient transformer Shen et al. (2021) to approximation the common transformer in order to reduce the memory and preserve the accuracy of calculating.
@@
3.6.4 Location masks
After getting two scores, we simply multiply two scores to get a final score. For every modality, it has one final score. This final score can comprehensively evaluate whether a point is suitable to add message. According to the number of message patches for one modality, we use top-k to select the most suitable locations for adding, where k is the number of patches. If we want to generate the next modality’s location mask, we should use the previous masks to avoid selecting the duplicate locations.
3.7 Learnable Selection Gate
This gate is to find out the locations where the information added. For above we can know that the self-adaption gate can produce a mask. Here we see the masks as the ground truth. After we get the ordered point set with added messages, we see it as the input of the learnable selection gate. The gate is a four layers MLP to predict the true location.
3.8 XD-Injection/Extraction Heads
These two are used for injecting messages and extracting messages. For each modal, we use a separate MLP Qi et al. (2017). Usually, 3 to 5 layers are enough for the task. We take out the GS points of target location recommended by self-adaption gate and just use the spherical harmonic parameters of three channels. We concatenate the spherical harmonic parameters with the message patches we want to added. We treat the concatenated data as input to multi-head injector which produce the spherical harmonic parameters with extra information.
@@ 3.8 XD-Injection/Extraction Heads
3.9 Training Process
Figure 3: This figure shows the training process of X-SG2S. Orange lines mean there exists gradient flow. Blue lines mean the gradient are truncated. Purple lines mean the GS cloud is sorted by sorting method. Black lines mean the losses used to train.
This model can be trained directly by GS cloud files or a pretrained 3DGS pipeline.
  1. 1.
@@ 3.9 Training Process
  2. 2.
    Use sorting method to sort the GS cloud to get an ordered set. Use the data loader to load the goal information we want to add. And we scatter them into message patches in order.
  3. 3.
@@
We can see from above that, in training steps we do not need to restore the message patches to their original structure. For example in our work, we encode a graph into features using AE. We inject the feature patches into GS points. If we want to restore the graph from AE, we just need to reshape the feature patches to the features and take it as the input of the decoder of the AE without optimize the encoder and decoder. And you do not need to design or use another losses to train. The message patches will automatically converge. And specifically for 3DGS small object watermarks, we should first reuses some of the points in the object to align the number of points in order to let X-SG2S perform parallel computation without causing extra loss.
3.10 Training Loss
@@
To ensure the learnable selection gate can precisely recognize the location where we inject the message, we use a mask cross entropy loss. This loss use the mask produced by self-adapted gate as the ground truth. The weight for the mask cross-entropy loss does not need to be set as its training is offline. We default this weight to 1.
@@
4 Experiments
4.1 Datasets
We use a pretrained model MVSplat for real time inference. We use large-scale ACID datasets to let MVSplatChen et al. (2025) generate the 3DGS scenes. For 1D dataset, we randomly generate a sequence of binary message. For each bit, it has the same probability to be 1 or 0. For 2D dataset, we use Logo-2KWang et al. (2020) dataset. For 3D dataset, we first download a subset of objaverseDeitke et al. (2023), and use GambaShen et al. (2024) to generate the 3DGS objects. They are then saved in .ply format as a 3D dataset.

### v2

@@
Abstract.
3D Gaussian Splatting (3DGS) has been widely used in 3D reconstruction and 3D generation. Training to get a 3DGS scene often takes a lot of time and resources and even valuable inspiration. The increasing amount of 3DGS digital asset have brought great challenges to the copyright protection. However, it still lacks profound exploration targeted at 3DGS. In this paper, we propose a new framework X-SG2S which can simultaneously inject 1 to 3D watermarks while keeping the original 3DGS scene almost unchanged. Generally, we have a X-SG2S injector for adding multi-modal messages simultaneously and an extractor for extract them. Specifically, we first split the watermarks into message patches in a fixed manner. A self-adaption gate is used to pick out suitable location for watermarking. Then use a XD(multi-dimension)-injection heads to add multi-modal messages into sorted 3DGS points. A learnable gate can recognize the location with extra messages and XD-extraction heads can restore hidden messages from the location recommended by the learnable gate. Extensive experiments demonstrated that the proposed X-SG2S can effectively conceal multi modal messages without changing pretrained 3DGS pipeline or the original form of 3DGS parameters. To more comprehensively evaluate the generalizability and practicality of our model, we assessed it from several aspects: the ability to maintain watermark and original 3DGS consistency, robustness against model degradation, and the model’s ability to make correct judgments. The experimental results demonstrate that our model performs exceptionally well in these areas. Meanwhile, with simple and efficient model structure and high practicality, X-SG2S still shows good performance in hiding and extracting multi-modal inner structured or unstructured messages. X-SG2S is the first to unify 1 to 3D watermarking model for 3DGS and the first framework to add multi-modal watermarks simultaneous in one 3DGS which pave the wave for later researches.
3DGS, Multi Modal Watermarks, Copyright Preservation
ccs: Security and privacy Social aspects of security and privacy
1. Introduction
Figure 1. Application scenario of the proposed. Scenario 1: after training a 3DGS pipeline by a trainer, an X-SG2S injector should be combined for joint use. In this way, when other users employ the trainer’s model, they can obtain a 3DGS result with some watermarks. By providing the 3DGS file to X-SG2S’s extractor, the trainer can determine whether a user has generated the result by using the trainer’s model. Scenario 2: if a user has specific requirements to add extra information to the 3DGS result for a particular purpose, they can directly use X-SG2S to achieve this goal. For example, you can add a text with “ACMMM”, an image of the ICML logo and a small 3DGS object at the same time by using X-SG2S’s injector and extract them by using X-SG2S’s extractor.
@@ 1. Introduction
Currently, only a small portion of research has addressed these problems. The new studies GS-Hider(zhang2024gs, ) and Splat in Splat(guo2024splats, ) restore image or 3D scene into another 3D scene. Considering the security issues of generative 3D, GaussianStego(li2024gaussianstego, ) explores how to embed image watermarks into diffusion models. GaussianMarker(huang2024gaussianmarker, ) embed 1D watermark into 3DGS. However, these models all have their limitations. For example, GS-Hider modifies the parameters of GS, which may increase the size of the 3D model parameters and no longer make it universal. Splat in Splat and GS-Hider can add watermarks only when they have the whole dataset of two scene. And they cannot add watermarks when user only have 3DGS model and watermarks message. GaussianMarker can only inject little binary message into 3DGS. The approach of GaussianStego rooting dino into diffusion can not well demonstrate its effect in feed forward 3D generative models. At the same time, adding extra watermarks in original 3DGS pipelines for protecting copyright is costly and complicate in design or training. We can see that how to effectively add additional watermarks without changing the GS parameters or model structure and let the method be general will be a great challenge. Meanwhile, embedding multimodal watermarks into a 3DGS simultaneously is even more of an unexplored area.
To solve the above challenges, we propose an effective, simple and flexible watermarking framework X-SG2S. It aims to embed 3D objects or images or binary messages simultaneously into the original 3DGS scene, and accurately extract the hidden message via just an ultra simple way. Specifically, we use patch the raw data and create redundancy which can let the watermark extracting process more robust. We innovatively proposed a model feature sparse DCAE for the restoration image when the some features are loss. Then, we use a self-adaption gate which can learn the adding position through the interaction between the additional watermarks and the 3DGS scenes. A XD-injection heads are used to add different modal messages respectively. Then a learnable gate is used to recognize the locations where we have added the messages. Finally, XD-extraction heads are used to restore the message added. The training for X-SG2S does not need to fine tune or redesign the pipeline. This property makes this model suitable not only for the feedforward model, but also for the model like diffusion which needs multiple iterations. And if you do not need to add additional information, you can use the original model in the fastest time without further fine-tuning. However, it is also possible to fine tune with the pipeline. Extensive experiments demonstrated that the proposed X-SG2S can effectively conceal multimodal messages without compromising rendering quality and possesses exceptional fidelity, robustness, veracity and flexibility. In a nutshell, the contributions and advantages of our X-SG2S can be summarized as follows.
@@
  • We create a model that can restore the whole image when some features are missing.
@@
2. Related Works
2.1. Generalizable GS
3D Gaussian Splatting (3DGS) (kerbl20233d, ) has emerged as a powerful method for reconstructing and representing 3D scenes using millions of 3D Gaussians. There has been a significant amount of work that has achieved good results in the reconstruction of small objects (szymanowicz2024splatter, ) (boss2024sf3d, ) and scenes (chen2025mvsplat, ) (charatan2024pixelsplat, ). Additionally, some studies have utilized diffusion models to realize text-to-3D generation (li2024instant3d, ), which also achieves promising outcomes.
2.2. 3D Steganography
Steganography has been evolving over the decades. Some research has made 3D steganography achieve good results in explicit geometry like meshes and point clouds (ohbuchi2002frequency, ) (zhu2024rethinking, ) (ferreira2020robust, ) or implicit geometer like nerf (li2023steganerf, )(luo2023copyrnerf, ). 3DGS is a new technology to represent explicit geometry. However, few works have been researched in steganography for 3DGS. GS-Hider (zhang2024gs, ) design a coupled secured feature attribute to replace the original 3DGS’s spherical harmonics coefficients and then use a scene decoder and a message decoder to disentangle the original RGB scene and the hidden message. Splat in Splat (guo2024splats, ) first train two scene by using the same location and shape of the cover scene and different color and opacity. Then they inject the secret scene into the cover one by an unlearnable method. And use an AE to map the transparency variation of each point. This method can work only when you have the entire multi view datasets. And for every scene, you should learn a new AE which is very fussy. GaussianStego(li2024gaussianstego, ) use dino to add image watermarks and use U-Net to extract them via multi rendered views. But there is not research about watermarking 1 to 3D messages to single 3DGS scene simultaneous while keeping GS parameters or model structure unchanged. GaussianMarker(huang2024gaussianmarker, ) can only embed 48 bit watermarks. The added information might be insufficient in specific contexts.
3. Safe Generalizable GS with X-dimensional watermarks
3.1. Preliminaries of Gaussian Splatting
@@ 3.2. Task Settings
3.3. Deep Thinking About Watermarking 3DGS
@@
  • If we want to know how to extract the information from the point set, we should firstly know which of them are added. This means you would better know the location of each point and the restoration order.
3.4. Overview
Figure 2. This figure shows the structure of self-adaption gate.
@@ 3.4. Overview
3.5. Patch the raw data and increase the number of patches
We need to patch the raw data because they usually have large size which can not be injected to single point. However, if we just inject the raw data patches, it will let the watermarks vulnerable to attack. If some of the GS points are hostile removed, we have no choice to restore all the data patches which means the watermarks will be destroyed. We need to find out a method to create patches from raw data, which can restore the whole watermark when the scene is attacked.
As shown in the Figure 4, for 1D and 3D data, we just duplicate the raw date many times because they are already patches. By repeating them, we can restore the watermarks by any subset of these data without difficulties. But for 2D message, things may be different. If we directly split the image into patches, we will get a large number of patches. At the same time, we cannot restore the whole image by any of the subset. So, we introduce a new method, feature sparse DCAE, which can restore the whole image by using subset of feature patches.
In the Figure 3, we show our feature sparse DCAE. Thanks for powerful ability of DCAE(chen2024deep, ), we can simply encode one 256 revolution image into a feature of shape $4\times 4\times 128$. Then, we flatten the feature and see them as a sequence. We use 4 up sampling convolution layers to create redundant context of this sequence and add fixed position encoding. In order to simulate the hostile attack, we ramdomly remove some of the patches. But this time, because we have a sequence with context, we have the ability to restore the whole sequence by any of the subset. Specifically, we using a MLP to predict the fixed position encoding, then use the predict encoding and the complete encoding to do nearest neighbor algorithm which find out where should the patches really located. We let all the patches be placed in the correct position. Then we use down sampling convolution layers and refinement MLP to restore origin feature sequence. Finally, we rearrange them and decode the watermark.
Figure 3. This figure shows the structure of feature sparse DCAE.
Figure 4. This figure shows we how to redundantize and patch the raw data to let the watermarking more robust.
3.6. Self-Adaption Gate
@@
This component is to find out the points which are suitable for injecting the messages while ensuring the original scenario is not disturbed so much. A simple idea is that let a gate to learn a score to measure the locations. The gate will give every point a comprehensive score. A higher score indicates that the location is more suitable for adding information.
@@
The comprehensive score is calculated from two scores:“self-score” and “cross-score”. “Self-score” is used to measure which points can significantly impact the 3DGS scenario and which are trivial. But only use “self-score” is not enough. We need to know where are more suitable for message to inject and extract. “Cross-score” is used to manage this problem.
@@
After we get two scores, we simply multiply two scores together to get the final score. By using top-k, we can set the recommended k positions to 1 and others to 0 to get the selection mask. The structure is shown in Figure 2.
3.6.1. message embedding
@@
Different modalities belong to different vector spaces. If we want to calculate the interactions results uniformly and efficiently, we need to map them into a unified space first. Here we simply use three different linear layer to map 1,2,3D and GS cloud message to a general space.
3.6.2. induced set attention to produce self-score
@@
After we get embedded message, we can measure the interact scores. Attention mechanism is a very efficient method for interact different information. We use transformer-based model to calculate the “self-score”. But we find that if we simply use scale dot product attention, the memory will overflow because the number of GS points are large. However, to calculate “self-score” is merely want to measure the importance of individual information in the overall information instead of learning detailed information between individuals. So set attention(lee2019set, ) is used to overcome the memory problem. At the same time, set transformer does not need for the data to be sequential. This is very suitable for point sets, because point sets are unordered.
@@
3.6.3. efficient attention to produce cross-score
Because “Cross-score” needs to interact the whole GS points with embedded message patches. However, if the amount of message patches is large, the attention may also face the problem of overflowing. Thus, we use efficient transformer (shen2021efficient, ) to approximation the common transformer in order to reduce the memory and preserve the accuracy of calculating.
@@
3.6.4. Location masks
After getting two scores, we simply multiply two scores to get a final score. For every modality, it has a final score. This final score can comprehensively evaluate whether a point is suitable to add message. We use top-k to select the most suitable locations for adding, where k is the number of patches. If we want to generate the next modality’s location mask, we should use the previous masks to avoid selecting the duplicate locations.
3.7. Learnable Selection Gate
For above we can know that the self-adaption gate can produce a mask. Here we see the masks as the ground truth. The learnable selection gate see ordered point set with added messages as the input to find out the locations where the information added. The gate is a four layers MLP to predict the true location.
3.8. Importance of SH
According to the Splat in Splat(guo2024splats, ), different order of SH has different importance. Higher-order SH contains only a small amount of High-frequency information and has less contribution to the scene. And we also do the experiments to prove this concept. So, in our machine, we choose to add watermark patches to order 3 and 4 of the SH.
3.9. XD-Injection/Extraction Heads
These two are used for injecting messages and extracting messages. For each modal, we use a separate MLP (qi2017pointnet, ). Usually, 3 to 5 layers are enough for the task. We take out the GS points of target location recommended by self-adaption gate and just use the spherical harmonic parameters of three channels. We concatenate the spherical harmonic parameters with the message patches we want to added. We treat the concatenated data as input to multi-head injector which produce the spherical harmonic parameters with extra information.
@@ 3.9. XD-Injection/Extraction Heads
3.10. Training Process
Figure 5. This figure shows the structure and training process of X-SG2S. Orange lines mean there exists gradient flow. Blue lines mean the gradient are truncated. Purple lines mean the GS cloud is sorted by sorting method. Black lines mean the losses used to train.
This model can be trained directly by GS cloud files or a pretrained 3DGS pipeline. In Figure 5, we show this process.
  1. (1)
@@ 3.10. Training Process
  2. (2)
    Preprocess the raw data into patches.
  3. (3)
@@
3.11. Training Loss
@@
To ensure the learnable selection gate can precisely recognize the location where we inject the message, we use a mask cross entropy loss. The weight for the mask cross-entropy loss does not need to be set as its training is offline. We default this weight to 1.
@@
4. Experiments
4.1. Datasets
We use a pretrained model MVSplat for real time inference. We use large-scale ACID datasets to let MVSplat(chen2025mvsplat, ) generate the 3DGS scenes. For 1D dataset, we randomly generate a sequence of binary message. For each bit, it has the same probability to be 1 or 0. For 2D dataset, we use Logo-2K(wang2020logo, ) dataset. For 3D dataset, we first download a subset of objaverse(deitke2023objaverse, ), and use Gamba(shen2024gamba, ) to generate the 3DGS objects. They are then saved in .ply format as a 3D dataset.